OpenStudy (anonymous):

Find the sum of the infinite series 4 + 3.2 + 2.56 + 2.048 +...if it exists.

OpenStudy (anonymous):

@Loser66 can u help?

OpenStudy (anonymous):

i cant remeber what the formula for sum is

OpenStudy (anonymous):

Sn=a1*1-n/1-n i think?

OpenStudy (welshfella):

this is a geometric series sum to infinity = 1 / (1 - r) where a = frist term and r = common difference

OpenStudy (welshfella):

sorry that' s a / (1 - r)

OpenStudy (welshfella):

to find r divide 2nd term by first this should equal 3rd / 2nd

OpenStudy (anonymous):

4/(r-.8)

OpenStudy (anonymous):

4/(1-.8)

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

i got .05

OpenStudy (welshfella):

you ve gone wrong somewhere = 4 / (1 - 0.8) = 4 / 0.2

OpenStudy (anonymous):

20

OpenStudy (welshfella):

thats it

OpenStudy (anonymous):

how would i know if it doesnt exists?

OpenStudy (welshfella):

here each term is decreasing - so the sum converges to a certain value If each term was greater than the last the sum would diverge - a limit would not exist

OpenStudy (anonymous):

okay. can u help me with a few more please? im still trying to get the hang out this

OpenStudy (welshfella):

eg 2 + 4 + 6 diverges

OpenStudy (welshfella):

ok

OpenStudy (anonymous):

should i ask here or open a new post

OpenStudy (welshfella):

best to open a new post